Can P(C) be found using the method shown in the photo?

  • Thread starter Thread starter desmond iking
  • Start date Start date
  • Tags Tags
    Probability
Join the discussion
Registration is free. Start your own thread to ask a follow-up.
3 replies · 2K views · last reply
desmond iking
Messages
284
Reaction score
2

Homework Statement



i have problem of finding P(C) , why can't i use the method as in the photo to find P(C) beacuse 1-(all unused stamps are missing)= at least 1 use stamps is missing.

Homework Equations





The Attempt at a Solution

 
Attachments
  • Printed probability exam question with a table of 24 stamps by colour (red, blue, yellow, green) and status (used, unused), followed by events A, B and C and parts a) and b).
    IMG_20140722_015315[1].jpg
    32.5 KB · Views: 461
  • Handwritten probability working on lined paper: P(A) = 20C3/24C3 = 285/506; P(B) = (21C1 x 22C2)/24C3 = 21/92; P(A and B) = (21C1 x 18C2)/24C3 = 153/1012; P(A or B) = 285/506 + 21/92 - 153/1012 = 162/253; and P(C) = 1 - (all unused stamps missing) = 1 - 14C3/24C3 = 91/506.
    DSC_0044[1].jpg
    27.4 KB · Views: 463
Physics news on Phys.org
C is the event at least one of the used stamps is missing.

"all unused stamps are missing" would mean there are 14 stamps missing - but the problem statement says only three are missing.

Did you mean, "all missing stamps are "unused"" - but that is not correct either since at least one of them has to be "used".

even C is the same as "at most two of the missing stamps are unused"

You need to comment your working.
 
Ya I mean all the 3 stamps missing are unused.
 
So if you write: D:=event all three missing stamps are unused, then you want to argue that P(C)=1-P(D), is that what you are saying?